Data for Hermann Corporation are shown below: Selling price Variable expenses Contribution margin Per Unit $ 90 63 Percent of2-a. Refer to the original data. How much will net operating income increase (decrease) per month if the company uses higher-Net operating income byShould the higher-quality components be used? Yes ONO
Answer #1
Current Proposed
Unit sales 2000 (2000*1.1)=2200
Sales (2000*90)=180,000 (2200*90)=198000
Variable expenses (2000*63)=126,000 2200*(63+2)=143,000
Contribution margin 54,000 55000
Fixed expenses 30000 30000
Net operating income 24000 25000

Hence Net operating income increases by =(25000-24000)=$1000

Hence higher quality components should be used(Yes)